Find exact solution:

|3-2x|=8
Thanks

Answer by stanbon(75887)

Find exact solution:
|3-2x|=8
3-2x = 8 or 3-2x = -8
2x = -5 or 2x = 11
x = -5/2 or x = 11/2
=========================
Cheers,
Stan H.